OpenStudy (clamin):

Please help! The second term of arithmetic sequence is 7. The sum of the first 4 terms of the arithmetic sequence is 12. Find the the first term a1, and tje common difference, d, of the sequence.

OpenStudy (welshfella):

2nd term = a1 + d where a1 = first term and d = common difference so we can write a + d = 7

OpenStudy (welshfella):

* a1 + d

OpenStudy (welshfella):

do you know the formula for the sum of n terms of an arithmetic sequence?

OpenStudy (welshfella):

hmm. i'm talking to myself...

OpenStudy (clamin):

an= a1 + (n -1)d

OpenStudy (mathmale):

Why not let n= {1, 2, 3, 4} and write out \[a _{n}=a _{1}+(n-1)d \]

OpenStudy (mathmale):

for each of those n values? This might help you find the unknowns, which are \[a _{1}~and~r\] ... where a1 is the first term and r is the common difference (not common ratio).

OpenStudy (mathmale):

Note that because the second term is 7, the formula given about would come out as follows for n=2 (2nd term):\[a _{2}=a _{1}+(2-1)d\]=7

OpenStudy (mathmale):

Thus, \[a _{2}=7=a _{1}+(2-1)d\]

OpenStudy (clamin):

What do i do nxt? To find the first term?

OpenStudy (anonymous):

This question needs attention. Its been long time I was dealing with these things

OpenStudy (mathmale):

What must you do to "solve" this problem? What are your goals? If you can answer these questions, then perhaps "what to do next?" will be clearer.

OpenStudy (clamin):

@welshfella